# Optimal design of unimorph-type cantilevered piezoelectric energy harvesters using level set-based topology optimization by considering manufacturability

### v1

@@ 4.1 Optimization algorithm
    The fictitious physical function field $p$ based on Eq. (57) is solved using the finite element method (FEM).
@@
    The eigenvector fields $\bm{u}_{oc}$ and $\bm{u}_{sc}$ and eigenvalues $\omega_{oc}$, $\omega_{sc}$, and $\varphi_{oc}$ derived from Eqs. (54)–(56) are solved using the FEM.
@@
    The objective functions $F_{pe}$ and $F_{sb}$ derived from Eq. (9) together with the output voltage $V_{E}$ at a frequency $\bar{\omega_{i}}$ and the constraint function $G_{V}$, which is based on Eq. (58), are evaluated.
@@
FreeFEM++ Hecht (2012) was used as the FEM solver. In the following subsections, we explain the approximations of the displacement and electrical potential fields and the sensitivity analysis.
@@
Within the Eulerian coordinate system, the fixed design domain necessitates the generation of finite elements during each iteration of the optimization process. To mitigate the computational expenditure, an ersatz material approach, as delineated in Allaire et al. (2004), was employed. In particular, the void domain was assumed to be a structural material with a relatively small elastic tensor, and the material properties were assumed to be smoothly distributed in the vicinity of the interface. The governing equations, which dictate the displacement field within the FEM, originally pertain to the material domain $\Omega_{D}$. Utilizing the level set, characteristic, and fictitious physical functions delineated in Section 3, coupled with an approximated Heaviside function $h(\phi)$ to be presented subsequently, we extended these equations to encompass the entire fixed design domain $D$ as follows:
$\begin{split}\chi(\phi)&\approx h(\phi),\\ h(\phi):&=\left\{\begin{array}[]{lll}d&\mathrm{for}\quad\phi<-w\\ \left\{\frac{1}{2}+\frac{\phi}{w}\left[\frac{15}{16}-\frac{\phi^{2}}{w^{2}}(\frac{5}{8}-\frac{3}{16}\frac{\phi^{2}}{w^{2}})\right]\right\}(1-d)+d&\mathrm{for}\quad-w\leq\phi\leq w\\ 1&\mathrm{for}\quad w<\phi\\ \end{array},\right.\end{split}$ (59)
${\bm{C}_{\phi ps}^{E}}:$ $=\bm{C}^{E}_{pe}\chi_{ps}h(\phi_{pe})h(p)+\bm{C}^{E}_{sb}\chi_{sp}h(\phi_{sb}),$ (60)
${\bm{e}}_{\phi ps}:$ $=\bm{e}\chi_{ps}h(\phi_{pe})h(p),$ (61)
${\bm{\varepsilon}}^{S}_{\phi ps}:$ $=\varepsilon_{0}\bm{I}(1-(\chi_{ps}h(\phi_{pe})h(p)+\chi_{sp}h(\phi_{sb})))+\bm{\varepsilon}^{S}\chi_{ps}h(\phi_{pe})h(p),$ (62)
${\rho}_{\phi ps}:$ $=\rho_{pe}\chi_{ps}h(\phi_{pe})h(p)+\rho_{sb}\chi_{sp}h(\phi_{sb}),$ (63)
@@ 4.3 Sensitivity analysis
Here, we describe a procedure for determining the sensitivity to update level set functions. In this study, based on Eqs. (54)–(58), we defined Lagrangian $L_{pe}$ and $L_{sb}$ with respect to $\phi_{pe}$ and $\phi_{sb}$, respectively, as follows:
$\begin{split}L_{pe}=&F_{pe}(\omega_{oc},\omega_{sc})+a_{oc}(\phi_{pe},\phi_{sb},\omega_{oc},\bm{u}_{oc},\bm{v}_{ocpe},\varphi_{oc})+b_{oc}(\phi_{pe},\phi_{sb},\bm{u}_{oc},\varphi_{oc},v_{\varphi_{ocpe}})\\ &+a_{sc}(\phi_{pe},\phi_{sb},\omega_{sc},\bm{u}_{sc},\bm{v}_{scpe})+a_{p}(\phi_{pe},\phi_{sb},p,v_{p})+\lambda G_{V}(\phi_{pe},\varphi_{oc}),\end{split}$ (64)
$\begin{split}L_{sb}=&F_{sb}(\omega_{oc},\omega_{sc})+a_{oc}(\phi_{pe},\phi_{sb},\omega_{oc},\bm{u}_{oc},\bm{v}_{ocsb},\varphi_{oc})+b_{oc}(\phi_{pe},\phi_{sb},\bm{u}_{oc},\varphi_{oc},v_{\varphi_{ocsb}})\\ &+a_{sc}(\phi_{pe},\phi_{sb},\omega_{sc},\bm{u}_{sc},\bm{v}_{scsb})+a_{p}(\phi_{pe},\phi_{sb},p,v_{p}),\end{split}$ (65)
where $\bm{v}_{ocpe}$, $\bm{v}_{ocsb}$, $\bm{v}_{sc}$, $\bm{v}_{scpe}$, $\bm{v}_{scsb}$, $v_{\varphi_{ocpe}}$, $v_{\varphi_{ocsb}}$, and $\lambda$ are the Lagrange multipliers. Using the adjoint method, sensitivity was determined as follows:
$\begin{split}{F}^{\prime}_{pe}=&c_{ocpe}\left\{a_{oc}(1,\phi_{sb},\omega_{oc},\bm{u}_{oc},\bm{u}_{oc},\varphi_{oc})+b_{oc}(1,\bm{u}_{oc},\varphi_{oc},\varphi_{oc})\right\}\\ &+c_{scpe}a_{sc}(1,\phi_{sb},\omega_{sc},\bm{u}_{sc},\bm{u}_{sc})+\lambda,\end{split}$ (66)
$\begin{split}{F}^{\prime}_{sb}=&c_{ocsb}\left\{a_{oc}(\phi_{pe},1,\omega_{oc},\bm{u}_{oc},\bm{u}_{oc},\varphi_{oc})\right\}\\ &+c_{scsb}\left\{a_{sc}(\phi_{pe},1,\omega_{sc},\bm{u}_{sc},\bm{u}_{sc})\right\}.\end{split}$ (67)
@@ 5 Numerical examples
The design domain and associated boundary conditions are illustrated in Figure 2. Let $D_{pe}$ denote a square, colored in blue in the figure, with a side length of 500 mm and a thickness of 4 mm. The object in this domain was a piezoelectric material. Similarly, let $D_{sb}$ symbolize a square, indicated in grey in the figure, with a side length of 500 mm and a thickness of 36 mm. The object within this domain did not exhibit any piezoelectric properties. The domain measuring 500 mm in length, 20 mm in width, and 36 mm in thickness, depicted in dark purple in the figure, was a nondesigned domain inhabited by a nonpiezoelectric material. Correspondingly, the domain that is 500 mm in length, 20 mm in breadth, and 4 mm in thickness, illustrated in light purple in the figure, represents a nondesigned domain occupied by a piezoelectric material. A fixed boundary condition was imposed on the surfaces adjacent to as well as the opposite domain $D_{sb}$. The domain colored red on the right side of the figure, delineating a 20-mm square with a thickness of 40 mm, is designated as the weight domain, which is a nondesigned domain filled with objects devoid of piezoelectric properties. The density was assumed to be 100 times that of the substrate material within this weight domain. The number of modes to be contemplated was fixed at $n=4$, and the specified natural frequencies were $[\bar{\omega}_{0},\bar{\omega}_{1},\bar{\omega}_{2},\bar{\omega}_{3}]=[70\;\text{Hz},435\;\text{Hz},450\;\text{Hz},500\;\text{Hz}]$. The initial structure was conjectured to be such that $\phi_{pe}=1$ in $D_{pe}$ and $\phi_{sb}=1$ in $D_{sb}$, indicating that the entire fixed design domain was occupied by the material domain.
@@ 5.1 Effectiveness of the proposed constraints for manufacturability
We conducted numerical experiments under nine conditions, as listed in Table 1. Each condition was characterized by a distinct setting of the regularization factor $\tau_{z}$ in the thickness direction, the representation of level set functions (either a singular function $\phi$ or a pair of functions $\phi_{pe}$ and $\phi_{sb}$), and the status of constraint implementation via a fictitious field $p$. Specifically, for each condition, the regularization factor $\tau_{z}$ was set to 1.0×$\times 10^{-6}$, 1.0×$\times 10^{-4}$, or 1.0×$\times 10^{-2}$. Furthermore, the level set functions were either represented solely by $\phi$ or by a combination of $\phi_{pe}$ and $\phi_{sb}$, and the constraints were applied using the fictitious field $p$ for some conditions. Each of these nine configurations was evaluated through 1000 iterations of computation. The parameters corresponding to each condition, coupled with the average values of the objective function $F_{k}$ and $F_{\omega}$ taken over the last 100 iterations, are also presented in Table 1. Under all conditions, $\alpha_{pe}$ and $\alpha_{sb}$ were set to $0.95$. The resulting structures for each condition are shown in Figure 3.
@@
The tradeoff relationship between the objective functions $F_{k}$ and $F_{\omega}$ for conditions (a)–(i) delineated in Table 1 is depicted in Figure 4. From the figure, a clear tradeoff relationship can be discerned between $F_{\omega}$ and $F_{k}$, the balance of which is influenced by changes in the parameter $\tau_{z}$. Moreover, conditions (d)–(f) yielded smaller $F_{k}$ values than conditions (a)–(c), given the same $\tau_{z}$ value. This phenomenon is attributed to the fact that, in conditions (d)–(f), the architecture of the piezoelectric film remained uninfluenced by the structure of the substrate, and, apart from regularization with respect to $\phi_{pe}$, it does not adhere to any constraints. This allows for a wider variety of potential solutions, thereby enhancing the structural degrees of freedom of piezoelectric films. In conditions (g)–(i), although the structure of the piezoelectric film was influenced by the structure of the substrate, the converse does not hold true, resulting in $F_{\omega}$ values that are lower than those under conditions (d)–(f).
@@
Table 1: Summary of parameter settings and objective functions: regularization factor $\tau_{z}$ in the thickness direction, the representation of level set functions, and the status of constraint implementation via a fictitious field $p$ and the corresponding objective functions $F_{k}$ and $F_{\omega}$ for each condition.
@@
condition $\tau_{z}$ level set function $p$ constraint $F_{k}$ $F_{\omega}$
@@
(d) 1.0×$\times 10^{-6}$ $\phi_{pe}\;$ and $\;\phi_{sb}$ not constrained 4.50×$\times 10^{-5}$ 1.51×$\times 10^{-5}$
(e) 1.0×$\times 10^{-4}$ $\phi_{pe}\;$ and $\;\phi_{sb}$ not constrained 5.80×$\times 10^{-5}$ 1.33×$\times 10^{-5}$
(f) 1.0×$\times 10^{-2}$ $\phi_{pe}\;$ and $\;\phi_{sb}$ not constrained 7.20×$\times 10^{-5}$ 1.08×$\times 10^{-5}$
(g) 1.0×$\times 10^{-6}$ $\phi_{pe}\;$ and $\;\phi_{sb}$ $p$ constrained 4.70×$\times 10^{-5}$ 1.25×$\times 10^{-5}$
(h) 1.0×$\times 10^{-4}$ $\phi_{pe}\;$ and $\;\phi_{sb}$ $p$ constrained 4.53×$\times 10^{-5}$ 1.28×$\times 10^{-5}$
(i) 1.0×$\times 10^{-2}$ $\phi_{pe}\;$ and $\;\phi_{sb}$ $p$ constrained 8.14×$\times 10^{-5}$ 4.79×$\times 10^{-6}$
Figure 3: Optimal configuration of an unimorph cantilevered energy harvester for conditions (a)–(i) listed in Table 1. The red region denotes the piezoelectric material domain $\Omega_{pe}$, and the blue region represents the substrate material domain $\Omega_{sb}$.
@@
For each condition, we computed two normalized quantities: $N_{\phi 1}$ and $N_{\phi 2}$. $N_{\phi 1}$ is the proportion of nodes at which the product of the level set functions with their adjacent nodes in the z-negative direction is negative, suggesting a boundary with the adjacent nodes. Conversely, $N_{\phi 2}$ is the proportion of nodes not located at the boundaries of domains $D_{sb}$ and $D_{pe}$. These values were normalized by the total number of nodes. The computed values for $N_{\phi 1}$ and $N_{\phi 2}$ for each condition are provided in Table 2.
As delineated in Table 2, irrespective of the level set function configuration and the existence of a $p$ constraint, an increase in the $\tau_{z}$ value corresponded to a reduction in the z-directional complexity. When two level set functions were utilized, as in conditions (d)–(i), the $N_{\phi 1}$ value surpassed that in conditions (a)–(c), where a single level set function embodied the entirety of the design domain. Contrastingly, $N_{\phi 2}$, excluding the interface between $D_{sb}$ and $D_{pe}$, did not exhibit a significant difference from $N_{\phi 1}$ under conditions (a)–(c) but was less than $N_{\phi 1}$ in conditions (d)–(i). This phenomenon can be attributed to the fact that, when two level set functions are employed, the structural complexity is not restrained at the boundary between domains $D_{sb}$ and $D_{pe}$, even upon altering the $\tau_{z}$ value. Hence, the $N_{\phi 1}$ value was considered higher than that in conditions (a)–(c).

### v2

@@ 4.1 Optimization algorithm
    The fictitious physical function field ${\xi}$ based on Eq. (101) is solved using the finite element method (FEM).
@@
    The eigenvector fields $\bm{u}_{oci}$ and $\bm{u}_{sci}$ and eigenvalues $\omega_{oci}$, $\omega_{sci}$, and $\varphi_{oci}$ derived from Eqs. (98)–(100) are solved using the FEM.
@@
    The displacement $\bm{u}$, the electric potential $\varphi_{n}$, the objective functions $F_{pe}$ and $F_{sb}$ derived from Eq. (23) together with the output voltage $V_{E}$ at a frequency $\bar{\omega_{i}}$ and the constraint function $G_{V}$, which is based on Eq. (102), are evaluated.
@@
All computations, including those involving the FEM solver, are carried out using FreeFEM++ [58]. In this study, convergence is considered achieved when the ratio of the objective function value from the previous iteration to that of the current iteration remains below $1.0\times 10^{-6}$ for 10 consecutive iterations. Additionally, the computation is terminated if the number of iterations reaches 1000. In the following subsections, we explain the approximations of the displacement and electrical potential fields and the sensitivity analysis.
@@
Within the Eulerian coordinate system, the fixed design domain necessitates the generation of finite elements during each iteration of the optimization process. To mitigate the computational expenditure, an ersatz material approach, as delineated in [56], is employed. In particular, the void domain is assumed to be a structural material with a relatively small Young’s modulus, and the material properties are assumed to be smoothly distributed in the vicinity of the interface between the weak and strong phases. The governing equations, which dictate the displacement field within the FEM, originally pertain to the material domain $\Omega$. Utilizing the level set, characteristic, and fictitious physical functions delineated in Section 3, coupled with an approximated Heaviside function $h(\phi)$ to be presented subsequently, we extend these equations to the entire fixed design domain $D$ as follows:
$\begin{split}\chi_{\phi}(\phi)&\approx h(\phi),\\ h(\phi):&=\left\{\begin{array}[]{lll}d&\mathrm{for}\quad\phi<-w\\ \left\{\frac{1}{2}+\frac{\phi}{w}\left[\frac{15}{16}-\frac{\phi^{2}}{w^{2}}(\frac{5}{8}-\frac{3}{16}\frac{\phi^{2}}{w^{2}})\right]\right\}(1-d)+d&\mathrm{for}\quad-w\leq\phi\leq w\\ 1&\mathrm{for}\quad w<\phi\\ \end{array},\right.\end{split}$ (103)
${\bm{C}_{\phi ps}^{E}}:$ $=\bm{C}^{E}_{pe}{h(\phi_{ps})}h({\phi_{p\psi}})h({\xi})+\bm{C}^{E}_{sb}{h(\phi_{sp})}h({\phi_{s\psi}}),$ (104)
${\bm{e}}_{\phi ps}:$ $=\bm{e}{h(\phi_{ps})}h({\phi_{p\psi}})h({\xi}),$ (105)
${\bm{\varepsilon}}^{S}_{\phi ps}:$ $=\varepsilon_{0}\bm{I}(1-({h(\phi_{ps})}h({\phi_{p\psi}})h({\xi})+{h(\phi_{sp})}h({\phi_{s\psi}})))+\bm{\varepsilon}^{S}{h(\phi_{ps})}h({\phi_{p\psi}})h({\xi}),$ (106)
${\rho}_{\phi ps}:$ $=\rho_{pe}{h(\phi_{ps})}h({\phi_{p\psi}})h({\xi})+\rho_{sb}{h(\phi_{sp})}h({\phi_{s\psi}}),$ (107)
@@ 4.3 Sensitivity analysis
Here, we describe a procedure for determining the sensitivity to update level set functions. In this study, based on Eqs. (98)–(102), we define the Lagrangian functions $L_{pe}$ and $L_{sb}$ of the level set functions with respect to ${\phi_{p\psi}}$ and ${\phi_{s\psi}}$, respectively, as follows:
$\begin{split}L_{pe}=&F_{pe}+{\sum_{i=1}^{n}\biggl{(}}a_{oci}({\phi_{p\psi}},{\phi_{s\psi}},\omega_{oci},\bm{u}_{oci},\bm{v}_{ocpe},\varphi_{oci})+b_{oci}({\phi_{p\psi}},{\phi_{s\psi}},\bm{u}_{oci},\varphi_{oci},v_{\varphi_{ocpe}})\\ &+a_{sci}({\phi_{p\psi}},{\phi_{s\psi}},\omega_{sci},\bm{u}_{sci},\bm{v}_{scpe}){\biggr{)}}+a_{\xi}({\phi_{p\psi}},{\phi_{s\psi}},{\xi},v_{\xi})+\lambda G_{V}({\phi_{p\psi}},\varphi_{n}),\end{split}$ (108)
$\begin{split}L_{sb}=&F_{sb}+{\sum_{i=1}^{n}\biggl{(}}a_{oci}({\phi_{p\psi}},{\phi_{s\psi}},\omega_{oci},\bm{u}_{oci},\bm{v}_{ocsb},\varphi_{oci})+b_{oci}({\phi_{p\psi}},{\phi_{s\psi}},\bm{u}_{oci},\varphi_{oci},v_{\varphi_{ocsb}})\\ &+a_{sci}({\phi_{p\psi}},{\phi_{s\psi}},\omega_{sci},\bm{u}_{sci},\bm{v}_{scsb}){\biggr{)}}+a_{\xi}({\phi_{p\psi}},{\phi_{s\psi}},{\xi},v_{\xi}),\end{split}$ (109)
where $\bm{v}_{ocpe}$, $\bm{v}_{ocsb}$, $\bm{v}_{scpe}$, $\bm{v}_{scsb}$, $v_{\varphi_{ocpe}}$, $v_{\varphi_{ocsb}}$, and $\lambda$ are the Lagrange multipliers. Using the adjoint method, sensitivity is determined as follows:
$\begin{split}{F}^{\prime}_{pe}=&{\sum_{i=1}^{n}\biggl{(}}c_{ocpe}\left\{a_{oci}(1,{\phi_{s\psi}},\omega_{oci},\bm{u}_{oci},\bm{u}_{oci},\varphi_{oci})+b_{oci}(1,\bm{u}_{oci},\varphi_{oci},\varphi_{oci})\right\}\\ &+c_{scpe}a_{sci}(1,{\phi_{s\psi}},\omega_{sci},\bm{u}_{sci},\bm{u}_{sci}){\biggr{)}}+\lambda,\end{split}$ (110)
$\begin{split}{F}^{\prime}_{sb}=&{\sum_{i=1}^{n}\biggl{(}}c_{ocsb}\left\{a_{oci}({\phi_{p\psi}},1,\omega_{oci},\bm{u}_{oci},\bm{u}_{oci},\varphi_{oci})\right\}\\ &+c_{scsb}\left\{a_{sci}({\phi_{p\psi}},1,\omega_{sci},\bm{u}_{sci},\bm{u}_{sci})\right\}{\biggr{)}}.\end{split}$ (111)
@@ 5 Numerical examples
The design domain and associated boundary conditions are illustrated in Figure 7. Let $D_{pe}$ denote a square, colored in blue in the figure, with a side length of 500 mm and a thickness of 4 mm. The material in this domain is PZT, which is a piezoelectric material. Similarly, let $D_{sb}$ symbolize a square, indicated in grey in the figure, with a side length of 500 mm and a thickness of 36 mm. The material within this domain is Silicon, which is a nonpiezoelectric material. The fixed design domain $D$ consists of the combined regions of $D_{pe}$ and $D_{sb}$. The domain measuring 500 mm in length, 20 mm in width, and 36 mm in thickness, depicted in dark purple in the figure, is a nondesign domain composed of Silicon. This domain will be referred to as the ”Nondesign domain (Silicon)”.
Correspondingly, the domain that is 500 mm in length, 20 mm in breadth, and 4 mm in thickness, illustrated in light purple in the figure, represents a nondesign domain filled with PZT. This domain will be referred to as the ”Nondesign domain (PZT)”.
A fixed boundary condition is imposed on the surface of the Nondesign domain (Silicon) that is opposite to the surface adjacent to the domain $D_{sb}$. The domain colored red on the right side of the figure, delineating a 20-mm square with a thickness of 40 mm, is designated as the weight domain, which is a nondesign domain filled with objects devoid of piezoelectric properties. To reduce computational load, the weight part is represented as a nondesign domain with a high density. The density is assumed to be 100 times that of the substrate material within this weight domain. This domain will be referred to as the ”Nondesign domain (Weight)”. The mesh configuration for different domains within this example is illustrated in Figure 8. In the nondesign domain (PZT), the mesh size is uniformly set to 10 $\times$ 10 $\times$ 2. The design domain $D_{pe}$ (PZT) also has a uniform mesh size of 20 $\times$ 10 $\times$ 2. For the nondesign domain (Silicon), the mesh size is 20 $\times$ 10 $\times$ 8.5 for $z$ coordinates less than 34, and 20 $\times$ 10 $\times$ 2 for $z$ coordinates 34 and above, near the interface with $D_{pe}$. The nondesign domain (Weight) follows the same pattern, with a mesh size of 20 $\times$ 10 $\times$ 8.5 for $z$ coordinates less than 34, and 20 $\times$ 10 $\times$ 2 for $z$ coordinates 34 and above. Finally, the design domain $D_{sb}$ (Silicon) has a mesh size of 20 $\times$ 10 $\times$ 8.5 for $z$ coordinates less than 34, except at the interface with $D_{pe}$, where the mesh size is 20 $\times$ 10 $\times$ 2 for $z$ coordinates 34 and above. For each condition, the computation time took approximately 100 hours. The number of modes to be contemplated is fixed at $n=4$, and the specified natural frequencies are $[\bar{\omega}_{0},\bar{\omega}_{1},\bar{\omega}_{2},\bar{\omega}_{3}]=[70\;\text{Hz},435\;\text{Hz},450\;\text{Hz},500\;\text{Hz}]$. The initial structure is conjectured to be such that ${\phi_{p\psi}}=1$ in $D_{pe}$ and ${\phi_{s\psi}}=1$ in $D_{sb}$, indicating that the entire fixed design domain is occupied by the material domain.
@@
Figure 8: Mesh configuration in different domains. In the nondesign domain (PZT), the mesh size is uniformly set to 10 $\times$ 10 $\times$ 2. The design domain $D_{pe}$ (PZT) has a uniform mesh size of 20 $\times$ 10 $\times$ 2. For the nondesign domain (Silicon), the mesh size is 20 $\times$ 10 $\times$ 8.5 for $z$ coordinates less than 34, and 20 $\times$ 10 $\times$ 2 for $z$ coordinates 34 and above, near the interface with $D_{pe}$. The nondesign domain (Weight) has a mesh size of 20 $\times$ 10 $\times$ 8.5 for $z$ coordinates less than 34, and 20 $\times$ 10 $\times$ 2 for $z$ coordinates 34 and above. The design domain $D_{sb}$ (Silicon) has a mesh size of 20 $\times$ 10 $\times$ 8.5 for $z$ coordinates less than 34, except at the interface with $D_{pe}$, where the mesh size is 20 $\times$ 10 $\times$ 2 for $z$ coordinates 34 and above.
@@ 5.1 Effectiveness of the proposed constraints for manufacturability
In this subsection, we present numerical examples to confirm the effectiveness of the constraints for manufacturability proposed in Section 3. In order to verify the effect of ”Constraint: the same cross-sectional shape throughout the design domain” mentioned in subsection 3.1.1, we test the regularization factor $\tau$. Similarly, to verify the effect of ”Constraint: the same cross-sectional shape in each domain” discussed in subsection 3.1.2, we examine the representation of level set functions. Lastly, to verify the effect of ”Constraint: no piezoelectric material is placed without a substrate” described in subsection 3.2, we evaluate the status of constraint implementation via a fictitious field ${\xi}$.
@@
We conduct numerical experiments under nine conditions, as listed in Table 1. Each condition is characterized by a distinct setting of the regularization factor $\tau_{z}$ in the thickness direction, the representation of level set functions either a singular function $\phi$ or a pair of functions ${\phi_{p\psi}}$ and ${\phi_{s\psi}}$, and the status of constraint implementation via a fictitious field ${\xi}$. Specifically, for each condition, the regularization factor $\tau_{z}$ is set to 1.0×$\times 10^{-6}$, 1.0×$\times 10^{-4}$, or 1.0×$\times 10^{-2}$. Furthermore, the level set functions are either represented solely by $\phi$ or by a combination of ${\phi_{p\psi}}$ and ${\phi_{s\psi}}$, and the constraints are applied using the fictitious field ${\xi}$ for conditions (g)–(i). Each of these nine configurations is evaluated through 1000 iterations of computation. The parameters corresponding to each condition, coupled with the average values of the objective function $F_{k}$ and $F_{\omega}$ taken over the last 100 iterations, are also presented in Table 1. Under all conditions, the weighting factors $\alpha_{pe}$ and $\alpha_{sb}$ are set to $0.95$ to achieve a balanced contribution between $F_{k}$ and $F_{\omega}$ in both the piezoelectric and the substrate domains. The resulting structures for each condition are shown in Figure 9.
@@
The tradeoff relationship between the objective functions $F_{k}$ and $F_{\omega}$ for conditions (a)–(i) delineated in Table 1 is depicted in Figure 10. From the figure, a clear tradeoff relationship can be discerned between $F_{\omega}$ and $F_{k}$, the balance of which is influenced by changes in the parameter $\tau_{z}$. Moreover, conditions (d)–(f) yield smaller $F_{k}$ values than conditions (a)–(c), given the same $\tau_{z}$ value. This phenomenon is attributed to the fact that, in conditions (d)–(f), the architecture of the piezoelectric film remains uninfluenced by the structure of the substrate, and, apart from regularization with respect to ${\phi_{p\psi}}$, it does not adhere to any constraints. This allows for a wider variety of potential solutions, thereby enhancing the structural degrees of freedom of piezoelectric films. In conditions (g)–(i), although the structure of the piezoelectric film is influenced by the structure of the substrate, the converse does not hold true, resulting in $F_{\omega}$ values that are lower than those under conditions (d)–(f).
@@
Table 1: Summary of parameter settings and objective functions: regularization factor $\tau_{z}$ in the thickness direction, the representation of level set functions, and the status of constraint implementation via a fictitious field ${\xi}$ and the corresponding objective functions $F_{k}$ and $F_{\omega}$ for each condition.
@@
condition $\tau_{z}$ level set function ${\xi}$ constraint $F_{k}$ $F_{\omega}$
@@
(d) 1.0×$\times 10^{-6}$ ${\phi_{p\psi}}\;$ and $\;{\phi_{s\psi}}$ not constrained 4.50×$\times 10^{-5}$ 1.51×$\times 10^{-5}$
(e) 1.0×$\times 10^{-4}$ ${\phi_{p\psi}}\;$ and $\;{\phi_{s\psi}}$ not constrained 5.80×$\times 10^{-5}$ 1.33×$\times 10^{-5}$
(f) 1.0×$\times 10^{-2}$ ${\phi_{p\psi}}\;$ and $\;{\phi_{s\psi}}$ not constrained 7.20×$\times 10^{-5}$ 1.08×$\times 10^{-5}$
(g) 1.0×$\times 10^{-6}$ ${\phi_{p\psi}}\;$ and $\;{\phi_{s\psi}}$ ${\xi}$ constrained 4.70×$\times 10^{-5}$ 1.25×$\times 10^{-5}$
(h) 1.0×$\times 10^{-4}$ ${\phi_{p\psi}}\;$ and $\;{\phi_{s\psi}}$ ${\xi}$ constrained 4.53×$\times 10^{-5}$ 1.28×$\times 10^{-5}$
(i) 1.0×$\times 10^{-2}$ ${\phi_{p\psi}}\;$ and $\;{\phi_{s\psi}}$ ${\xi}$ constrained 8.14×$\times 10^{-5}$ 4.79×$\times 10^{-6}$
Figure 9: Optimal configuration of a unimorph cantilevered energy harvester for conditions (a)–(i) listed in Table 1. The red region denotes the piezoelectric material domain $\Omega_{pe}$, and the blue region represents the substrate material domain $\Omega_{sb}$.
@@
Next, the following paragraph explains the effectiveness of the proposed constraints. We compute two normalized quantities: $N_{\phi 1}$ and $N_{\phi 2}$. $N_{\phi 1}$ is calculated to confirm the extent to which boundaries in the $z$ direction exist within the entire design domain $D$. This $N_{\phi 1}$ verifies whether the constant cross-sectional constraint is effective throughout the fixed design domain $D$. $N_{\phi 1}$ is the proportion of nodes where the level set function changes sign with adjacent nodes in the negative $z$-direction, indicating a boundary. Conversely, $N_{\phi 2}$ is the proportion of nodes , among those extracted by $N_{\phi 1}$, that are not located at the boundaries of domains $D_{sb}$ and $D_{pe}$. By calculating $N_{\phi 2}$, we verify whether the constraint of having a constant cross-section in each of $D_{pe}$ and $D_{sb}$ is effective. These values are normalized by the total number of nodes. The computed values for $N_{\phi 1}$ and $N_{\phi 2}$ for each condition are provided in Table 2.
As delineated in Table 2, irrespective of the level set function configuration and the existence of a ${\xi}$ constraint, an increase in the $\tau_{z}$ value corresponds to a reduction in the z-directional complexity. When two level set functions are utilized, as in conditions (d)–(i), the $N_{\phi 1}$ value surpasses that in conditions (a)–(c), where a single level set function embodied the entirety of the design domain. Contrastingly, $N_{\phi 2}$, excluding the interface between $D_{sb}$ and $D_{pe}$, does not exhibit a significant difference from $N_{\phi 1}$ under conditions (a)–(c) but is less than $N_{\phi 1}$ in conditions (d)–(i). This phenomenon can be attributed to the fact that, when two level set functions are employed, the structural complexity is not restrained at the boundary between domains $D_{sb}$ and $D_{pe}$, even upon altering the $\tau_{z}$ value. Hence, the $N_{\phi 1}$ value is considered higher than that in conditions (a)–(c).